OpenStudy (anonymous):

Help with some rational equation. y-2/y=y/y-1. Will rewrite with the equation thing.

OpenStudy (anonymous):

\[\frac{ y-2 }{ y}=\frac{ y }{ y-1 }\]

OpenStudy (anonymous):

Start by cross multiplying- \[y^{2} = (y-2)(y-1) = y^{2} -3y +2\] Subtract y^2 from both sides: \[0=2-3y\] Add 3y to both sides: \[3y=2\] so \[y=\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

cross multiplying y-2by y-1 and y by y correct? Thanks!

OpenStudy (anonymous):

how come we got rid of \[y^{2} \] like that? could you explain?

OpenStudy (anonymous):

We had y^2 on both sides, so we can subtract it from both sides.

OpenStudy (anonymous):

Thanks @vinnv226

OpenStudy (anonymous):

No problem, glad I could help.